OpenStudy (anonymous):

Just for fun: How do you get the answer 24 by only using the numbers 8,8,3,3 and the arithmetic operators?

OpenStudy (anonymous):

you can use (not necessarily all) +, -, /,* , ^ and parentheses.

OpenStudy (anonymous):

is ^ arithmetic operator?

OpenStudy (anonymous):

no i'm not sure that it is - its not needed anyway

OpenStudy (radar):

I can get 25, but so far not 24! lol

OpenStudy (anonymous):

close - i must admit this one took me about 30 minutes before i got it - and a headache!

OpenStudy (radar):

I made a copy of the question, and will continue wrestling with it.

OpenStudy (radar):

I am assuming you can only use the digits once.

OpenStudy (anonymous):

yes

OpenStudy (radar):

O.K. Hope your headache goes away as I think I got one coming on lol.

OpenStudy (anonymous):

\[8*(8/3+1/3)=24\]

OpenStudy (radar):

The number 1 is not on the list of allowed digits.

OpenStudy (anonymous):

errr...ok?

OpenStudy (radar):

Here I thought my headache was over lol

OpenStudy (cwrw238):

mathmuse you helped me - i got it 8 / (3 - 8/3) = 8 / 1.3 = 24

OpenStudy (anonymous):

nice one

OpenStudy (radar):

Ah yes congrats to cwrw238 and mathmuse. Now the headaches is gone!